OpenStudy (anonymous):

x+y+z=7100 .07x+.10y+.12z=759 -x-y+z=1100

OpenStudy (anonymous):

which method

OpenStudy (anonymous):

x+y+z=7100 -x-y+z=1100 =2z=8200 z=4100

OpenStudy (anonymous):

system of equation

OpenStudy (anonymous):

z is correct now try to solve for x and y, will help u when u get stuck

OpenStudy (anonymous):

x+y+4100=7100-4100= x+y=3000

OpenStudy (anonymous):

.07x+.10y+.12(4100)=759 ? then thats where i get stuck

OpenStudy (anonymous):

.07x+.10y+.12z=759 multiply by 10 then add -x-y+z=1100

OpenStudy (anonymous):

come on amanda u can do it

OpenStudy (anonymous):

@Mandy86 did u sleep

OpenStudy (anonymous):

yeah but ive been at this all day!

OpenStudy (anonymous):

thats y im here for:p .07x+.10y+.12z=759 multiply everything by 10 what do you get

OpenStudy (anonymous):

Thank you I dont know what i would do with you!

OpenStudy (anonymous):

you are welcome; now do wat i told u

OpenStudy (anonymous):

.7x+y+1.2z=7590

OpenStudy (anonymous):

.7x+y+1.2z=7590 -x-y+z=1100 add both equation

OpenStudy (anonymous):

great thank you!

OpenStudy (anonymous):

6x+2.2z=8690

OpenStudy (anonymous):

oh -.3x

OpenStudy (anonymous):

wanna double check it before i tell you what you did wrong

OpenStudy (anonymous):

-.3x+2.2z=8690

OpenStudy (anonymous):

what is z?

OpenStudy (anonymous):

2.2?

OpenStudy (anonymous):

you solved for it earlier

OpenStudy (anonymous):

4100

OpenStudy (anonymous):

so i plug it 4100 back in to z

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-.3x+2.2(4100)=8690

OpenStudy (anonymous):

now solve for x

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

x=1100

OpenStudy (anonymous):

no i plug in x and z in the first equation?

OpenStudy (anonymous):

plug in into any equation...the one you love the most lol

OpenStudy (anonymous):

to get z

OpenStudy (anonymous):

to get y, u have z already

OpenStudy (anonymous):

yeah i meant that!

OpenStudy (anonymous):

but u didnt write it :p

OpenStudy (anonymous):

i know! long day1

OpenStudy (anonymous):

hehe how was ur day

OpenStudy (anonymous):

long lol and stessful so much homework to do. Hows yours?

OpenStudy (anonymous):

y=1900

OpenStudy (anonymous):

yay! thank you where are you from?

OpenStudy (anonymous):

correct, ohio, u?

OpenStudy (anonymous):

Florida!

OpenStudy (anonymous):

what are you studying?

OpenStudy (anonymous):

College algebra and biology

OpenStudy (anonymous):

im in my 4th year of aa degree

OpenStudy (anonymous):

u graduating in may?

OpenStudy (anonymous):

yes then in fall i start my BA in elementary education!

OpenStudy (anonymous):

thats good; i will help u in math if u want, as long as u r trying and willing to learn

OpenStudy (anonymous):

I am i enjoying math its just it gets overwhelming.